# Homogeneous solutions to the Einstein-matter equations with a magnetic field and a conformal gauge singularity

### v1

@@ 4 The Einstein-Vlasov system with a magnetic field
$\begin{split}\mathcal{M}^{pqrs}&=\frac{2}{a}\Big{(}b^{pq}h^{r}h^{s}+2b^{rs}h^{p}h^{q}+\frac{1}{2}(a_{mn}h^{m}h^{n})(b^{pr}b^{qs}+b^{ps}b^{qr}-2b^{pq}b^{rs})\\ &\quad{}-b^{qr}h^{p}h^{s}-b^{pr}h^{q}h^{s}-b^{qs}h^{p}h^{r}-b^{ps}h^{q}h^{r}\Big{)}.\end{split}$ (89)
@@
where $G_{ij}$ should not be confused with the Einstein tensor.
@@
and the differentiability condition that $F$, $G$, $H$, $N$, $\frac{\partial F}{\partial x}$, $\frac{\partial F}{\partial y}$, $\frac{\partial G}{\partial x}$, $\frac{\partial G}{\partial y}$, $\frac{\partial H}{\partial\tau}$, $\frac{\partial H}{\partial x}$, $\frac{\partial^{2}H}{\partial\tau\partial x}$, $\frac{\partial^{2}H}{\partial x^{2}}$, $\frac{\partial N}{\partial\tau}$, $\frac{\partial N}{\partial x}$, $\frac{\partial^{2}N}{\partial\tau\partial x}$, $\frac{\partial^{2}N}{\partial x^{2}}$ should be continuous. Moreover we need that all the eigenvalues of $N(0,x_{0})$ have positive real part. If these conditions are satisfied there is a unique solution to the initial value problem (91) for all initial data satisfying the compatibility condition (102) and this solution depends continuously on $F$, $G$, $H$, $N$, $x_{0}$ and $y_{0}$. The solution is continuous for $\tau\geq 0$ and is continuously differentiable and satisfies the differential equation equation for $\tau\geq 0$. If all the eigenvalues of $N$ have real parts which are greater than $1$ then the solution is continuously differentiable and satisfies the equation for $\tau\geq 0$. We will obtain all these conditions. The last condition we will obtain by restricting the size of the norm of the magnetic field.
@@ 4.2 Fuchsian condition at $\tau=0$
is finite and positive. The particular symmetric bilinear form ${\delta^{ij}}$ appears to play a distinguished role here but it is easy to see that it could be replaced here by any other choice of positive definite symmetric bilinear form. We obtain
@@
Under our hypotheses on ${h}$ this minimum exists and is positive so
@@
The covariant Hessian of ${4b^{-1/2}\Psi}$ is computed as follows:
$4b^{-1/2}\Psi=16\pi\int\left(b^{mn}p_{m}p_{n}\right)^{1/2}f\,d^{3}p,$ (123)
$\begin{split}\nabla_{kl}\left(4b^{-1/2}\Psi\right)&=\delta_{kl}\left(4b^{-1/2}\Psi\right)=16\pi\int\delta_{kl}\left(b^{mn}p_{m}p_{n}\right)^{1/2}f\,d^{3}p\\ &=8\pi\int\left(b^{mn}p_{m}p_{n}\right)^{-1/2}p_{k}p_{l}f\,d^{3}p=2b^{-1/2}\Psi_{kl},\end{split}$ (124)
$\begin{split}\nabla_{ij}\nabla_{kl}\left(4b^{-1/2}\Psi\right)&=\delta_{ij}\nabla_{kl}\left(4b^{-1/2}\Psi\right)-\Gamma^{pq}_{ij,kl}\nabla_{pq}\left(4b^{-1/2}\Psi\right)\\ &=-4\pi\int\left(b^{mn}p_{m}p_{n}\right)^{-3/2}p_{i}p_{j}p_{k}p_{l}f\,d^{3}p\\ &\quad{}+\frac{1}{2}b^{-1/2}\left(a_{ik}\Psi_{jl}+a_{il}\Psi_{jk}+a_{jk}\Psi_{il}+a_{jl}\Psi_{ik}\right)\\ &=\frac{1}{2}b^{-1/2}\left(-2\Psi_{ijkl}+a_{ik}\Psi_{jl}+a_{il}\Psi_{jk}+a_{jk}\Psi_{il}+a_{jl}\Psi_{ik}\right).\end{split}$ (125)
@@
$\left\{\left[b^{mn}p_{m}p_{n}\left(a_{ik}p_{j}p_{k}+a_{il}p_{j}p_{k}+a_{jk}p_{i}p_{k}+a_{jl}p_{i}p_{k}\right)\right]-2p_{i}p_{j}p_{k}p_{l}\right\}c^{ij}c^{jk},$
@@
Finally, we have to consider the last term in ${F}$. We have that
$\begin{split}\nabla_{kl}\left(4b^{1/2}a_{mn}h^{m}h^{n}\right)&=\nabla_{kl}\left(4b^{1/2}\right)a_{mn}h^{m}h^{n}+b^{1/2}a_{mn}\nabla_{kl}\left(4h^{m}h^{n}\right)\\ &=2b^{1/2}\left(a_{kl}a_{mn}-a_{km}a_{ln}-a_{kn}a_{lm}\right)h^{m}h^{n},\end{split}$ (129)
from which follows that
@@
has exactly three critical points, at ${(0,\sqrt{1/2})}$, ${(0,-\sqrt{1/2})}$ and ${(2,0)}$. The first two of these are the only ones in the unit disc, which is a contractible set, and both are minima. Since the argument given here applies to the case ${h=0}$ it can serve as a replacement for the one given there.
@@
Let us consider two vectors $e^{i}$, $f^{i}$ which are orthogonal two each other and to the magnetic field, such that:
@@ 4.3 The eigenvalues of $N$ at $\tau=0$
We follow [4] and consider the eigenvalues of $\chi$ which we name $\mu$. Working in the basis where $a_{ij}=\delta_{ij}$ and $k_{ij}=\operatorname{diag}(k_{1},k_{2},k_{3})$ we obtain
$\mu|k_{i}|=\chi^{pq}_{ij}k^{ij}=4\pi\left|\int\left(\delta^{kl}p_{k}p_{l}\right)^{-\frac{3}{2}}p_{i}^{2}(p_{1}^{2}k_{1}+p_{2}^{2}k_{2}+p_{3}^{2}k_{3})f\,d^{3}p\right|$
@@
$\mu\sum_{i=1}^{3}|k_{i}|$ $\leq 4\pi\int\left(\delta^{kl}p_{k}p_{l}\right)^{-\frac{1}{2}}(p_{1}^{2}|k_{1}|+p_{2}^{2}|k_{2}|+p_{3}^{2}|k_{3}|)f=\Psi_{11}|k_{1}|+\Psi_{22}|k_{2}|+\Psi_{33}|k_{3}|$
@@
$\hat{k}_{ij}-b^{mr}b^{ns}\Psi_{ijrs}\hat{k}_{mn}+\Psi_{ij}\hat{k}-\mathcal{M}_{ij}^{mn}\hat{k}_{mn}=(\lambda-1)\left(\hat{k}_{ij}+\frac{1}{2}a_{ij}\hat{k}-\frac{1}{2}\lambda\hat{k}_{ij}\right),$ (145)
Contract with $b^{ij}$
@@
Since both $\chi^{mn}_{ij}$ and $S_{ij}^{mn}$ are symmetric $\nu$ will be real and bounded by the sum of the eigenvalues of $\chi^{mn}_{ij}$ and the eigenvalues of $S_{ij}^{mn}$, which implies using (139) and (160) that
@@
4.4 The main theorems of Einstein-Vlasov
Recall that the unknowns for the Einstein-Vlasov system with a magnetic field are $3\times 3$ real symmetric matrices, a vector and a distribution function. We denote by $S_{2}({\mathbb{R}}^{3})$ the space of $3\times 3$ symmetric matrices. We are now ready to conclude with the following theorem:
@@ 4.4 The main theorems of Einstein-Vlasov
Let $a_{0},b_{0},k_{0},{Z}_{0}\in S_{2}({\mathbb{R}}^{3})$, $h\in{\mathbb{R}}^{3}$, and $0\leq f_{0}$, a smooth function with compact support in $\mathbb{R}^{3}\setminus\{0\}$ which is not identically zero, be initial data of the rescaled Einstein-Vlasov system with a magnetic field with Bianchi I symmetry (76), (79)–(82) satisfying the Fuchsian conditions (103), (131) and the constraints (70), (75). Then, there exists a time interval $[0,T]$ on which the rescaled Einstein-Vlasov system has a unique solution $a_{ij},b^{ij},k_{ij},{Z}_{ij}\in C^{0}([0,T];S_{2}({\mathbb{R}}^{3}))$ and $f\in C^{1}([0,T];L^{1}({\mathbb{R}}^{3}))$. If in addition $(a_{ij}h^{i}h^{j})_{0}<\frac{1}{6}$, then we even have a unique differentiable solution $a_{ij},b^{ij},k_{ij},{Z}_{ij}\in C^{1}([0,T];S_{2}({\mathbb{R}}^{3}))$.
@@
However due to (131) $a_{1ij}=a_{2ij}=0$. We thus have established the following:
Theorem 2
$\tilde{a}_{ij}$ $=2ta_{0ij}+o(t^{\frac{3}{2}}),$ (178)
$\tilde{k}_{ij}$ $=2a_{0ij}+o(t^{\frac{1}{2}}).$ (179)
$\tilde{a}_{ij}$ $=2ta_{0ij}+o(t^{2}),$ (180)
$\tilde{k}_{ij}$ $=2a_{0ij}+o(t).$ (181)
@@ 5 The Einstein-Boltzmann system with a magnetic field
which is not regular at $\tau=0$. We replace $\tau$ by $s$ where $ds=\tau^{\gamma-2}d\tau$ so that
@@
and note that $\gamma-1>0$ by assumption. Later it will we useful to use the notation
@@
$\hat{Z}_{ij}=\tau^{2-\gamma}Z_{ij}=\frac{c_{\gamma}}{s}\left(\frac{8\pi}{\sqrt{a}}\int_{N_{x}}fp_{i}p_{j}\frac{d^{3}p}{(b^{mn}p_{m}p_{n})^{1/2}}+\frac{2}{a}(a_{ij}a_{mn}-2a_{im}a_{jn})h^{m}h^{n}-a_{ij}\right),$ (191)
@@
The differentiability conditions were shown in [15] and we already have seen that in the Vlasov case they are satisfied for the magnetic term. We now procede to compute the eigenvalues of the new $N$ which we again call $\lambda$. The eigentensors of $N$ satisfy
@@
$K_{ij}-b^{mr}b^{ns}\Psi_{ijrs}K_{mn}+\Psi_{ij}\hat{k}-\mathcal{M}_{ij}^{mn}K_{mn}=\left(\frac{\lambda}{c_{\gamma}}-1\right)\left(\frac{\gamma}{2}K_{ij}+\frac{1}{2}a_{ij}K-\frac{1}{2}\frac{\lambda}{c_{\gamma}}K_{ij}\right),$ (217)
@@
We can now translate this to the physical version using the results of [15] obtaining
Theorem 4
with
with
$\mathcal{C}_{ij}=\frac{1}{(\gamma-1)^{2}}2^{\gamma}a_{2ij},$ (236)
where $a_{2ij}$ is a constant which depends only on $f_{0}$ and $h$.
Acknowledgements
The authors acknowledge the hospitality from the Oberwolfach Research Institute for Mathematics where during the research stay under the Oberwolfach Research Fellow (OWRF) program with reference number 2427q the present work was finished in an excellent working environment.

### v2

@@ 4 The Einstein-Vlasov system with a magnetic field
$\begin{split}\mathcal{M}^{pqrs}&=\frac{2}{a}\Big{(}b^{pq}h^{r}h^{s}+2b^{rs}h^{p}h^{q}+\frac{1}{2}(a_{mn}h^{m}h^{n})(b^{pr}b^{qs}+b^{ps}b^{qr}-2b^{pq}b^{rs})\\ &\quad{}-b^{qr}h^{p}h^{s}-b^{pr}h^{q}h^{s}-b^{qs}h^{p}h^{r}-b^{ps}h^{q}h^{r}\Big{)},\end{split}$ (100)
@@
where $G_{ij}$ should not be confused with the Einstein tensor and where we have used the notation $M_{ij}^{mn}=a_{ip}a_{jq}M^{pqmn}$.
@@
and the differentiability condition that $F$, $G$, $H$, $N$, $\frac{\partial F}{\partial x}$, $\frac{\partial F}{\partial y}$, $\frac{\partial G}{\partial x}$, $\frac{\partial G}{\partial y}$, $\frac{\partial H}{\partial\tau}$, $\frac{\partial H}{\partial x}$, $\frac{\partial^{2}H}{\partial\tau\partial x}$, $\frac{\partial^{2}H}{\partial x^{2}}$, $\frac{\partial N}{\partial\tau}$, $\frac{\partial N}{\partial x}$, $\frac{\partial^{2}N}{\partial\tau\partial x}$, $\frac{\partial^{2}N}{\partial x^{2}}$ should be continuous. Moreover we need that all the eigenvalues of $N(0,x_{0})$ have positive real part. If these conditions are satisfied there is a unique solution to the initial value problem (102) for all initial data satisfying the compatibility condition (113) and this solution depends continuously on $F$, $G$, $H$, $N$, $x_{0}$ and $y_{0}$. The solution is continuous for $\tau\geq 0$ and is continuously differentiable and satisfies the differential equation equation for $\tau>0$. If all the eigenvalues of $N$ have real parts which are greater than $1$ then the solution is continuously differentiable and satisfies the equation for $\tau\geq 0$. We will obtain all these conditions. The last condition we will obtain by restricting the size of the norm of the magnetic field.
@@ 4.2 Fuchsian condition at $\tau=0$
is finite and positive. The particular symmetric bilinear form ${\delta^{ij}}$ appears to play a distinguished role here but it is easy to see that it could be replaced here by any other choice of positive definite symmetric bilinear form. In order to compute the covariant derivative of $F$ we observe that
$4b^{-1/2}\Psi=16\pi\int\left(b^{mn}p_{m}p_{n}\right)^{1/2}f\,d^{3}p,$ (124)
so that
$\nabla_{kl}\left(4b^{-1/2}\Psi\right)$ $=\partial_{kl}\left(4b^{-1/2}\Psi\right)=16\pi\int\partial_{kl}\left(b^{mn}p_{m}p_{n}\right)^{1/2}f\,d^{3}p$
$=8\pi\int\left(b^{mn}p_{m}p_{n}\right)^{-1/2}p_{k}p_{l}f\,d^{3}p=2b^{-1/2}\Psi_{kl},$ (125)
and
$\nabla_{kl}\left(4b^{1/2}a_{mn}h^{m}h^{n}\right)$ $=\nabla_{kl}\left(4b^{1/2}\right)a_{mn}h^{m}h^{n}+b^{1/2}a_{mn}\nabla_{kl}\left(4h^{m}h^{n}\right)$
$=2b^{1/2}\left(a_{kl}a_{mn}-a_{km}a_{ln}-a_{kn}a_{lm}\right)h^{m}h^{n}.$ (126)
We thus obtain
@@
Under our hypotheses on ${f}$ this minimum exists and is positive so
@@
The covariant Hessian of ${4b^{-1/2}\Psi}$ is computed using (125) as follows:
$\begin{split}\nabla_{ij}\nabla_{kl}\left(4b^{-1/2}\Psi\right)&=\partial_{ij}\nabla_{kl}\left(4b^{-1/2}\Psi\right)-\Gamma^{pq}_{ij,kl}\nabla_{pq}\left(4b^{-1/2}\Psi\right)\\ &=-4\pi\int\left(b^{mn}p_{m}p_{n}\right)^{-3/2}p_{i}p_{j}p_{k}p_{l}f\,d^{3}p\\ &\quad{}+\frac{1}{2}b^{-1/2}\left(a_{ik}\Psi_{jl}+a_{il}\Psi_{jk}+a_{jk}\Psi_{il}+a_{jl}\Psi_{ik}\right)\\ &=\frac{1}{2}b^{-1/2}\left(-2\Psi_{ijkl}+a_{ik}\Psi_{jl}+a_{il}\Psi_{jk}+a_{jk}\Psi_{il}+a_{jl}\Psi_{ik}\right).\end{split}$ (137)
@@
$\left\{\left[b^{mn}p_{m}p_{n}\left(a_{ik}p_{j}p_{l}+a_{il}p_{j}p_{k}+a_{jk}p_{i}p_{l}+a_{jl}p_{i}p_{k}\right)\right]-2p_{i}p_{j}p_{k}p_{l}\right\}c^{ij}c^{kl},$
@@
Finally, we have to consider the last term in ${F}$. Using (126) we obtain
@@
has exactly three critical points, at ${(0,\sqrt{1/2})}$, ${(0,-\sqrt{1/2})}$ and ${(2,0)}$. The first two of these are the only ones in the unit disc, which is a contractible set, and both are minima. So their argument is incorrect, but the argument given here applies to the case ${h=0}$ and it can serve as a replacement for the one given there.
@@
Let us consider two vectors $e^{i}$, $f^{i}$ which are orthogonal to each other and to the magnetic field, such that:
@@ 4.3 The eigenvalues of $N$ at $\tau=0$
We follow [4] and consider the eigenvalues of $\chi$ which we name $\mu$, i.e. $\chi^{pq}_{ij}k^{ij}=\mu k^{pq}$. Working in the basis where $a_{ij}=\delta_{ij}$ and $k_{ij}=\operatorname{diag}(k_{1},k_{2},k_{3})$ we obtain
$|\mu k_{i}|=\left|4\pi\int\left(\delta^{kl}p_{k}p_{l}\right)^{-\frac{3}{2}}p_{i}^{2}(p_{1}^{2}k_{1}+p_{2}^{2}k_{2}+p_{3}^{2}k_{3})f\,d^{3}p\right|$
@@
$\sum_{i=1}^{3}|\mu k_{i}|$ $\leq 4\pi\int\left(\delta^{kl}p_{k}p_{l}\right)^{-\frac{1}{2}}(p_{1}^{2}|k_{1}|+p_{2}^{2}|k_{2}|+p_{3}^{2}|k_{3}|)f=\Psi_{11}|k_{1}|+\Psi_{22}|k_{2}|+\Psi_{33}|k_{3}|$
@@
which using the definitions of $\chi^{mn}_{ij}$, $\Pi_{ij}^{mn}$ and $\pi_{ij}^{mn}$ (94) turns into
$\hat{k}_{ij}-b^{mr}b^{ns}\Psi_{ijrs}\hat{k}_{mn}+\Psi_{ij}b^{mn}\hat{k}_{mn}-\mathcal{M}_{ij}^{mn}\hat{k}_{mn}=(\lambda-1)\left(\hat{k}_{ij}+\frac{1}{2}a_{ij}b^{mn}\hat{k}_{mn}-\frac{1}{2}\lambda\hat{k}_{ij}\right),$
and using the notation $\hat{k}=b^{mn}\hat{k}_{mn}$:
$\hat{k}_{ij}-b^{mr}b^{ns}\Psi_{ijrs}\hat{k}_{mn}+\Psi_{ij}\hat{k}-\mathcal{M}_{ij}^{mn}\hat{k}_{mn}=(\lambda-1)\left(\hat{k}_{ij}+\frac{1}{2}a_{ij}\hat{k}-\frac{1}{2}\lambda\hat{k}_{ij}\right).$ (156)
If one contracts the last equation with $b^{ij}$ and uses (114), (143)-(144) one obtains
@@
Since both $\chi^{mn}_{ij}$ and $S_{ij}^{mn}$ are symmetric $\nu$ will be real and bounded by the largest eigenvalue of $\chi^{mn}_{ij}$ plus the largest eigenvalue of $S_{ij}^{mn}$, which implies using (150) and (171) that
@@
4.4 The main results of Einstein-Vlasov
Recall that the unknowns for the Einstein-Vlasov system with a magnetic field are $3\times 3$ real symmetric matrices, a vector and a distribution function. We are now ready to conclude with the following theorem:
@@ 4.4 The main results of Einstein-Vlasov
Let $a_{0},b_{0},k_{0},{Z}_{0}\in S_{2}({\mathbb{R}}^{3})$, $h\in{\mathbb{R}}^{3}$, and $f_{0}\geq 0$, a smooth function with compact support in $\mathbb{R}^{3}\setminus\{0\}$ which is not identically zero, be initial data of the rescaled Einstein-Vlasov system with a magnetic field with Bianchi I symmetry (87), (90)–(93) satisfying the Fuchsian conditions (114), (142) and the constraints (70), (76). Then, there exists a time interval $[0,T]$ on which the rescaled Einstein-Vlasov system has a unique solution $a_{ij},b^{ij},k_{ij},{Z}_{ij}\in C^{0}([0,T];S_{2}({\mathbb{R}}^{3}))$ and $f\in C^{1}([0,T];L^{1}({\mathbb{R}}^{3}))$. If in addition $(a_{ij}h^{i}h^{j})_{0}<\frac{1}{6}$, then we even have a unique differentiable solution $a_{ij},b^{ij},k_{ij},{Z}_{ij}\in C^{1}([0,T];S_{2}({\mathbb{R}}^{3}))$.
@@
However due to (142) $a_{1ij}=a_{2ij}=0$. We thus have established Theorem 1.
@@ 5 The Einstein-Boltzmann system with a magnetic field
which is not regular at $\tau=0$ and where $h$ denotes the unphysical relative momentum which should not be confused with the magnetic field. We replace $\tau$ by $s$ where $ds=\tau^{\gamma-2}d\tau$ so that
@@
and note that $\gamma-1>0$ by assumption. Later it will be useful to use the notation
@@
$\hat{Z}_{ij}=\tau^{2-\gamma}Z_{ij}=\frac{c_{\gamma}}{s}\left(\frac{8\pi}{\sqrt{a}}\int fp_{i}p_{j}\frac{d^{3}p}{(b^{mn}p_{m}p_{n})^{1/2}}+\frac{2}{a}(a_{ij}a_{mn}-2a_{im}a_{jn})h^{m}h^{n}-a_{ij}\right),$ (198)
@@
The differentiability conditions were shown in [15] and we already have seen that in the Vlasov case they are satisfied for the magnetic term. We now proceed to compute the eigenvalues of the new $N$ which we again call $\lambda$. The eigentensors of $N$ satisfy
@@
$K_{ij}-b^{mr}b^{ns}\Psi_{ijrs}K_{mn}+\Psi_{ij}\hat{k}-\mathcal{M}_{ij}^{mn}K_{mn}=\left(\frac{\lambda}{c_{\gamma}}-1\right)\left(\frac{\gamma}{2}K_{ij}+\frac{1}{2}a_{ij}K-\frac{1}{2}\frac{\lambda}{c_{\gamma}}K_{ij}\right).$ (224)
@@
We can now translate this to the physical version using the results of [15] and obtain Theorem 2.
Conflicts of interest/Competing interests
The authors acknowledge the hospitality from the Oberwolfach Research Institute for Mathematics where during the research stay under the Oberwolfach Research Fellow (OWRF) program with reference number 2427q the present work was finished in an excellent working environment. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00451692). Otherwise, the authors have no conflicts of interest to declare that are relevant to the content of this article.